OpenStudy (anonymous):

write an equation of the line perpendicular to the given line that contains P. P(4,4);y= -2x - 8

OpenStudy (anonymous):

Can you see what the slope of the line that it is perpendicular to is?

OpenStudy (anonymous):

no because this is all it says

OpenStudy (anonymous):

Well the equation of the line is y=-2x-8. Thats in the form y=mx+b where m=slope. So the slope is -2. A line perpendicular to that would have the negative reciprocal slope, 1/2. now that you know a point and the slope put it in y-y1=m(x-x1) where (x1,y1) is the point and m is the slope